OpenStudy (anonymous):

factor x2 + 3x + 9/4

OpenStudy (anonymous):

same process as well.

OpenStudy (anonymous):

example: x^2+5x+25/4=(x+5/2)(x+5/2)

OpenStudy (anonymous):

since, 5/2+5/2=5, and 5/2*5/2=25/4.

OpenStudy (theviper):

\[\Huge{\implies x^2+\frac{3x}{2}+\frac{3x}{2}+\frac{9}{4}}\]\[\LARGE{\implies x \left( x+\frac{3}{2} \right)}+\frac{3}{2}\left( x+\frac{3}{2} \right)=??\]

OpenStudy (theviper):

Understood???